OpenStudy (anonymous):

-4x^5y^-1/5 over 2x^-1/5y....Simplify and express the answer with positive exponents. Evaluate numerical expressions.

OpenStudy (anonymous):

okay...few parts, so bear with me

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[\frac{-4x^5y^{-1/5}}{2x^{-1/5}y}\]is this your question?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

Hello lokisan?

OpenStudy (anonymous):

\[\frac{-4x^5y^{-1/5}}{2x^{-1/5}y}=\frac{-4}{2}.\frac{x^5}{x^{-1/5}}.\frac{y^{-1/5}}{y}\]\[=-2.x^{5+1/5}.y^{-1/5-1}=-2x^{26/5}y^{-6/5}\]\[=-2\frac{x^{26/5}}{y^{6/5}}=-2\left( \frac{x^{13}}{y^3} \right)^{2/5}\]

OpenStudy (anonymous):

Do you understand the steps?

OpenStudy (anonymous):

I do, I got to the last part....But Im still a little confused on my answer part. I always get it wrong

OpenStudy (anonymous):

If I had more time, I'd try and go through it some more. What I can do is recommend this site: www.khanacademy.org - it's free. Look up "exponents" and "simplifying expressions with exponents". It may help - these clips are quite good. Good luck - just learn your rules for exponents and take everything step-by-step, don't try skipping anything until you're more experienced.

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

no probs.